OpenStudy (anonymous):

What is the greater of two numbers whose product is 700, if the sum of the two numbers exceeds their difference by 20?

OpenStudy (anonymous):

tips?

OpenStudy (anonymous):

how could I set this one up?

OpenStudy (anonymous):

I have no idea, let me check.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

x+x-20 = 700?

OpenStudy (anonymous):

70+10?

OpenStudy (anonymous):

70 times 10?

OpenStudy (anonymous):

70+10=80

OpenStudy (anonymous):

70-10=60

OpenStudy (anonymous):

70 is the greater of the two numbers

OpenStudy (anonymous):

thats what I got

OpenStudy (anonymous):

Haha, I can't help you with this one, sorry.

OpenStudy (anonymous):

70 is the anser

OpenStudy (anonymous):

answer

OpenStudy (anonymous):

Well, I have no idea how it could be.

OpenStudy (anonymous):

ty though

OpenStudy (anonymous):

70 times 10=700

OpenStudy (anonymous):

70+10=80

OpenStudy (anonymous):

70-10=60

OpenStudy (anonymous):

80-60=20

OpenStudy (anonymous):

70 times 10=700

OpenStudy (anonymous):

70 is the greatest integer

OpenStudy (anonymous):

so 70 is the answer

OpenStudy (anonymous):

ty though

OpenStudy (anonymous):

Sure.